the illuminance E at a point on a surface is given by the formula
E = I/dē
where E= illuminance in lux
I= luminous intensity of the light source in candelas
d= distance in metres
rearrange this formula to make d the subject and find the value of d given that E= 25 lux and I= 100 candelas

Answer by richwmiller(17219)

d^2=I/E
d^2=100/25
d^2=4
d=2 meters